### v1

@@ Abstract
Distributed Quantum Computing (DQC) is essential for scaling quantum algorithms beyond the limitations of monolithic NISQ devices. However, the current software ecosystem forces developers to manually orchestrate low-level network resources, such as entanglement generation (EPR pairs) and classical synchronization, leading to verbose, error-prone, and non-scalable code. This paper introduces NetQMPI, a high-level Python framework that adapts the Message Passing Interface (MPI) standard to the quantum domain using the Single Program Multiple Data (SPMD) paradigm. Built as a middleware over the NetQASM SDK, NetQMPI abstracts the underlying physical topology, automating network initialization and resource management through a unified Communicator interface. We propose semantic point-to-point primitives and novel collective operations—such as expose and unexpose—that address the constraints of the No-Cloning Theorem by leveraging multipartite entanglement for data distribution. Our comparative analysis demonstrates that NetQMPI decouples algorithmic logic from network size, reducing the code complexity for generating an $N$-node GHZ state from $\mathcal{O}(N^{2})$ to constant complexity $\mathcal{O}(1)$. Furthermore, the framework ensures backend agnosticism, enabling the seamless execution of high-level applications on rigorous physical simulators like NetSquid (via SquidASM) and future quantum hardware adhering to the NetQASM standard.
@@
distributed quantum computing, high-performance computing, quantum internet, mpi, quantum networks
@@ I Introduction
In recent years, quantum computing has emerged as a computational paradigm capable of offering exponential speedups for problems intractable on classical machines [Buhrman1998QuantumComputation, Markov2014LimitsComputation], such as integer factorization [Shor1999Polynomial-timeComputer] and unstructured search [Grover1996ASearch]. However, realizing these algorithms at scale requires a substantial number of qubits and robust error correction, capabilities that remain challenging for current hardware in the noisy intermediate-scale quantum (NISQ) era [Preskill2018QuantumBeyond].
To overcome these scaling limitations, distributed quantum computing (DQC) has emerged as a promising path. By interconnecting multiple quantum processing units (QPUs) via a quantum network, it is possible to create a virtual cluster capable of executing large-scale algorithms [Barral2025ReviewComputing, Caleffi2024DistributedSurvey, VanMeter2008ArchitectureAlgorithm]. However, orchestrating computations across such a distributed infrastructure introduces engineering complexities [Ovide2023MappingArchitectures]. Unlike classical networks, quantum networks require the management of fragile resources such as entanglement, Einstein-Podolsky-Rosen (EPR) pairs, and the tight synchronization of classical and quantum control planes [Wehner2018QuantumAhead].
Nevertheless, the transition from monolithic to distributed quantum algorithms introduces significant software engineering challenges [Loke2022FromOverview]. Most existing software tools allow to researchers to program at the physical or link layer, managing elementary operations explicitly [Cardama2026NetQIR:Computing, Diaz-Camacho2025BenchmarkingEmulators]. This low-level approach not only increases development complexity but also couples the algorithmic logic to specific hardware topologies or simulation backends, severely limiting code portability and maintainability. In the realm of classical high-performance computing (HPC), this challenge was successfully addressed by the message passing interface (MPI) standard, which unified distributed programming under the single program multiple data (SPMD) paradigm. While theoretical proposals for a quantum message passing interface (QMPI) exist [Haner2021DistributedQmpi], the ecosystem still lacks a functional, standardized implementation that bridges the gap between high-level algorithmic design and the execution on rigorous quantum network stacks.
@@ I-A Contributions
  • Adoption of the SPMD Paradigm: We propose a shift from role-based scripting to a unified Single Program Multiple Data model. By injecting logical rank identifiers, NetQMPI allows a single source file to orchestrate complex protocols across $N$ nodes, decoupling the code complexity from the network size.
  • Abstraction of Physical Resources: We introduce the QMPICommunicator and semantic primitives (e.g., qsend, qrecv) that abstract the management of EPR sockets and classical control planes. This allows researchers to focus on algorithmic logic rather than the intricate details of entanglement generation and fidelity management.
@@
This paper is organized as follows, Section II provides the necessary background on quantum networking tools through the different layers of abstraction, also explaining NetQASM SDK. Section III details the architecture of NetQMPI as well as the creation process on NetQASM SDK and the high-level semantic functions provided.. Section IV a comparative analysis of the various tools in the literature cited in the related work, and Section V concludes the paper.
@@ II-A Software Tools for Programming Distributed Quantum Systems
Interlinq [Parekh2021QuantumComputing], offers a high-level abstraction layer built upon the QuNetSim communication framework, designed to automate circuit partitioning and message scheduling across distributed nodes.
  • Strengths: It simplifies distributed execution by automating circuit partitioning and teleportation scheduling, allowing monolithic circuits to be run across multiple nodes (“circuit knitting”) without manual intervention for small-scale systems.
QuNetSim [Diadamo2021QuNetSim:Networks], created by the Technical University of Munich, provides a high-level Python API for simulating network protocols (e.g., quantum key distribution (QKD), teleportation) without dealing with hardware physics.
@@
QMPI [Haner2021DistributedQmpi], proposed by Häner et al. (ETH Zurich), introduces the theoretical foundation for applying the Message Passing Interface (MPI) standard to quantum computing.
@@ II-B Quantum Network Simulators
  • Integration: While it acts as a standalone distributed backend, it is part of the QuTech ecosystem and was a precursor to modern execution environments.
@@
  • Integration: SquidASM is explicitly built on top of NetSquid. It translates the high-level instructions from the SDK into discrete events that NetSquid can process, providing a realistic noise model while maintaining a programmable interface.
@@ II-C Discrete-Event Simulators
  • Strengths: Excellent for studying network topology and protocol stacks and available as open-source.
@@ II-D NetQASM SDK Overview
II-D1 NetQASM ISA
NetQASM defines a low-level, platform-independent Instruction Set Architecture designed specifically for quantum networks. It serves as a common language that abstracts the heterogeneity of underlying hardware. Programs compiled into NetQASM assembly can be executed on various backends, ranging from high-fidelity physical simulators like SquidASM (based on NetSquid) or SimulaQron. This architecture separates the quantum logic from the control hardware, allowing for hybrid quantum-classical execution.
@@ II-D1 NetQASM ISA
II-D2 The SDK and EPR Sockets
To facilitate programming without writing raw assembly, the NetQASM SDK provides a Python interface to write applications for specific network roles (e.g., “Alice” and “Bob”). A central abstraction in the SDK is the EPR Socket. By opening an EPR socket between two nodes, the programmer can request the generation of entanglement (e.g., creating Bell pairs) utilizing the underlying network link layer.
Figure LABEL:fig:netqasm-cnot demonstrates the implementation of a teleport protocol using the NetQASM SDK, highlighting the steps required to transfer a qubit state from Alice to Bob.
II-D3 Programming Complexity
While the SDK abstracts the physical generation of entanglement, it does not provide high-level primitives for qubit transmission. Crucially, to transfer a quantum state from one node to another, the generation of an EPR pair is merely the first step. The programmer must manually implement the full Quantum Teleportation or Remote State Preparation (Teledata) protocol ad hoc.
For example, to send a single qubit via teleportation using the SDK, the developer must explicitly:
@@ II-D3 Programming Complexity
    Request entanglement via the EPR Socket and wait for completion.
@@
III NetQMPI proposal
This section introduces NetQMPI, a high-level Python library designed to simplify the development of distributed quantum applications. Inspired by the standardized Message Passing Interface (MPI) used in classical High-Performance Computing (HPC), NetQMPI adopts the Single Program Multiple Data (SPMD) paradigm. This allows developers to write a single code, platform-independent source code that orchestrates quantum operations across multiple network nodes without managing the underlying physical entanglement resources manually.
III-A Software stack
The architectural position of NetQMPI within the quantum network software ecosystem is illustrated in Figure 3. NetQMPI acts as a middleware layer that sits directly on top of the NetQASM SDK.
Figure 3: The NetQMPI software stack. The library wraps the verbose NetQASM SDK primitives, providing a unified API for the application layer while relying on the standard NetQASM ISA to interface with backend simulators like SquidASM or SimulaQron.
@@ III-A Software stack
Because these primitives compile down to the standardized NetQASM Instruction Set Architecture (ISA), NetQMPI is inherently backend-agnostic. This abstraction allows developers to write code that is completely decoupled from the underlying infrastructure, enabling the execution of the same source code on both simulated networks and physical quantum networks without modification.
@@
III-B Construction of NetQMPI from NetQASM SDK
@@
The core orchestration is handled by the external.py module, which manages the multiprocess environment mapping. The network state and automated entanglement generation are encapsulated within the communicator package. Finally, the specific communication logic is organized under primitives, divided into p2p and collective modules, whose specific implementations will be detailed in Subsections III-C and III-D, respectively.
@@
III-B1 Single Program Architecture
In the standard NetQASM SDK, the developer is required to write $N$ separate source files (e.g., app_alice.py, app_bob.py), one for each node in the network. NetQMPI unifies this into a single file following the SPMD paradigm.
To enable this, NetQMPI utilizes the abstraction of the rank. For readers unfamiliar with HPC terminologies, the rank is simply a unique integer identifier assigned to each process in the distributed network, ranging from 0 to size1size-1. This value is injected automatically by the library into the user’s code at runtime, allowing the developer to define the logic for all nodes in a single function signature.
@@ 6 COMM_WORLD = QMPICommunicator(rank, size, app_config)
Figure 5: Example of the rank and size variable injection with the global communicator creation.
When the user executes this script using the Command Line Interface (CLI), the cli.py module parses the -n argument (number of processes) and triggers the external.py dispatcher. This dispatcher spawns the simulation processes using the NetQASM SDK, passing the specific rank to each instance. For example, executing the code with 3 nodes results in:
@@
As illustrated in Figure 6, the execution flow is orchestrated by external.py. It parses the unified script, defines the network roles based on the requested size, and injects the corresponding rank information into the entry point of the application. Finally, it hands over the configuration to the underlying NetQASM SDK, which manages the physical spawning of the runtime processes (e.g., Alice, Bob, Charlie).
@@
III-B2 Automated Network Initialization
The core abstraction that enables this connectivity is the QMPICommunicator class, defined in the communicator package.
Conceptually, a communicator acts as a virtualization layer that groups a set of logical processes connected through a logical network. This construct abstracts the underlying complexity of the infrastructure, where logical processes (identified by ranks) are mapped to physical resources (Quantum Processing Units) via a physical interconnection network. This relationships is visually described in Figure 7. Figure 7(a) depicts the physical reality where QPUs are connected via a Quantum Router, while Figure 7(b) illustrates how the software stack (OS/Middleware) maps the high-level Logical Ranks ($1\dots n$) to these physical resources transparently.
Figure 7: Logical abstraction for scheduling physical resources. (a) Example of physical resources interconnected between a physical quantum network. (b) The Communicator provides a logical layer where abstract Ranks are assigned to processes, and the operative system automatically handles the mapping to the available physical QPUs.
This approach mirrors the architecture of classical distributed computing, allowing developers to focus on the algorithmic logic rather than the physical topology. For a comprehensive definition of this layered abstraction model in distributed quantum computing tools, we refer the reader to the architectural framework and specifically Figure 6 presented in [Cardama2026NetQIR:Computing].
From an implementation perspective, the communicator automates the resource management that is manually handled in the raw SDK. In standard NetQASM, establishing a network requires the user to instantiate a NetQASMConnection and define specific EPRSocket objects for every pair of communicating nodes, leading to an $O(N^{2})$ initialization complexity in fully connected topologies. NetQMPI resolves this by performing the following automated steps during initialization:

### v2

@@ Abstract
Distributed Quantum Computing (DQC) is essential for scaling quantum algorithms beyond the limitations of monolithic NISQ devices. However, the current software ecosystem forces developers to manually orchestrate low-level network resources, such as entanglement generation (EPR pairs) and classical synchronization, leading to verbose, error-prone, and non-scalable code. This paper introduces NetQMPI, a high-level Python framework that adapts the Message Passing Interface (MPI) standard to the quantum domain using the Single Program Multiple Data (SPMD) paradigm. Built as a middleware over the NetQASM SDK, NetQMPI abstracts the underlying physical topology, automating network initialization and resource management through a unified Communicator interface. We propose semantic point-to-point primitives and novel collective operations—such as expose and unexpose—that address the constraints of the No-Cloning Theorem by leveraging multipartite entanglement for data distribution. Our comparative analysis demonstrates that NetQMPI decouples algorithmic logic from network size, reducing the code complexity for generating an $N$-node GHZ state from $\mathcal{O}(N^{2})$ to constant complexity $\mathcal{O}(1)$. Furthermore, the framework ensures backend agnosticism, enabling the seamless execution of high-level applications on rigorous physical simulators, such as NetSquid (via SquidASM), and future quantum hardware adhering to the NetQASM standard.
@@
Distributed Quantum Computing, High-Performance Computing, MPI, Quantum Networks.
@@ I Introduction
In recent years, quantum computing has emerged as a computational paradigm capable of offering exponential speedups for problems intractable on classical machines [Buhrman1998QuantumComputation, Markov2014LimitsComputation], such as integer factorization [Shor1999Polynomial-timeComputer] and unstructured search [Grover1996ASearch]. However, realizing these algorithms at scale requires a substantial number of qubits and robust error correction capabilities, which remain challenging for current hardware in the Noisy Intermediate-Scale Quantum (NISQ) era [Preskill2018QuantumBeyond].
To overcome these scaling limitations, Distributed Quantum Computing (DQC) has emerged as a promising path. By interconnecting multiple Quantum Processing Units (QPUs) via a quantum network, it is possible to create a virtual cluster capable of executing large-scale algorithms [Barral2025ReviewComputing, Caleffi2024DistributedSurvey, VanMeter2008ArchitectureAlgorithm]. However, orchestrating computations across such a distributed infrastructure introduces engineering complexities [Ovide2023MappingArchitectures]. Unlike classical networks, quantum networks require the management of fragile resources such as entanglement, Einstein-Podolsky-Rosen (EPR) pairs, and the tight synchronization of classical and quantum control planes [Wehner2018QuantumAhead, Illiano2022QuantumSurvey, Azuma2023QuantumInternet].
Nevertheless, the transition from monolithic to distributed quantum algorithms introduces significant software engineering challenges [Loke2022FromOverview]. Most existing software tools require researchers to program at the physical or link layer, where they must explicitly manage elementary operations [Cardama2026NetQIR:Computing, Diaz-Camacho2025BenchmarkingEmulators, Barral2025ReviewComputing]. At this level, developers are often burdened with defining hardware-specific parameters such as photon emission timing, fiber attenuation, and memory decoherence , while also manually orchestrating network protocols like entanglement generation (EPR pairs), routing, and the tight synchronization of classical and quantum control planes [Cacciapuoti2020QuantumComputing].
This low-level approach not only increases development complexity but also couples the algorithmic logic to specific hardware topologies or simulation backends, severely limiting code portability and maintainability. In the realm of classical High-Performance Computing (HPC), this challenge was successfully addressed by the Message Passing Interface (MPI) standard, which unified distributed programming, typically following the Single Program Multiple Data (SPMD) paradigm. While theoretical proposals for a Quantum Message Passing Interface (QMPI) exist [Haner2021DistributedQmpi], the ecosystem still lacks a functional, standardized implementation that bridges the gap between high-level algorithmic design and the execution on rigorous quantum network stacks.
@@ I-A Contributions
  • Adoption of the SPMD Paradigm: We propose a shift from role-based scripting to a unified Single Program Multiple Data model. By injecting logical rank identifiers, NetQMPI enables a single source file to orchestrate complex protocols across $N$ nodes, thereby decoupling code complexity from network size.
  • Abstraction of Physical Resources: We introduce the QMPICommunicator and semantic primitives (e.g., qsend, qrecv) that abstract the management of EPR sockets and classical control planes, which allows researchers to focus on algorithmic logic rather than the intricate details of entanglement generation and fidelity management.
@@
This paper is organized as follows. Section II introduces the related work including tools for DQC programming and simulation. Section III provides an overview of the NetQASM framework, distinguishing between its ISA and the SDK, while highlighting the programming complexity inherent in the current ecosystem. Section IV details the architecture of NetQMPI as well as the creation process on NetQASM SDK and the high-level semantic functions provided. Section V presents a comparative analysis of the various tools in the literature described in the related work section, and, finally, Section VI concludes the paper.
@@ II-A Software Tools for Programming Distributed Quantum Systems
QuNetSim [Diadamo2021QuNetSim:Networks] provides a high-level Python API for simulating network protocols (e.g., Quantum Key Distribution (QKD), teleportation) without dealing with hardware physics.
@@
Interlinq [Parekh2021QuantumComputing] offers a high-level abstraction layer built upon the QuNetSim communication framework, designed to automate circuit partitioning and message scheduling across distributed nodes.
  • Strengths: It simplifies distributed execution by automating circuit partitioning and teleportation scheduling, allowing monolithic circuits to be run across multiple nodes (“circuit knitting”, i.e., splitting a quantum program to fit onto smaller devices linked by classical communication [Piveteau2024CircuitCommunication]) without manual intervention for small-scale systems.
QMPI, proposed by Häner et al. [Haner2021DistributedQmpi], introduces the theoretical foundation for applying the Message Passing Interface (MPI) standard to quantum computing.
@@ II-B Quantum Network Simulators
  • Integration: While it acts as a standalone distributed backend, it is part of the QuTech ecosystem (a suite of standardized quantum networking tools developed at TU Delft) and was a precursor to modern execution environments.
@@
  • Integration: SquidASM is explicitly built on top of NetSquid (see next subsection). It translates the high-level instructions from the SDK into discrete events that NetSquid can process, providing a realistic noise model while maintaining a programmable interface.
@@ II-C Discrete-Event Simulators
  • Strengths: Excellent for studying network topology and protocol stacks, and available as open-source.
@@ III NetQASM SDK Overview
III-A NetQASM ISA
NetQASM defines a low-level, platform-independent ISA explicitly designed for quantum networks. It serves as a common language that abstracts the heterogeneity of underlying hardware. Programs compiled into NetQASM assembly can be executed on various backends, ranging from high-fidelity physical simulators like SquidASM (based on NetSquid) or SimulaQron. This architecture separates the quantum logic from the control hardware, allowing for hybrid quantum-classical execution.
To illustrate the granularity of control required by this architecture, Figure 1 presents a snippet of Alice’s operations. The execution flow begins with explicit resource management, where the program manually assigns a value to a classical register (set R0 0) to identify, allocate (qalloc), and initialize (init) a local control qubit. Subsequently, the entanglement generation process requires preparing the arguments in specific memory addresses—storing the number of pairs at @0 and the type at @1—before triggering the create_epr instruction. Notably, the code must include a wait_all instruction to explicitly block execution until the physical hardware confirms the entanglement generation. Finally, the script manually maps the remote entangled qubit to a register (set R1 1), performs a local two-qubit gate (cnot), measures the result into a classical register (meas), and explicitly frees the quantum memory with qfree.
@@ III-A NetQASM ISA
Alice (Control Node)
1# Create a socket to send classical information
2socket = Socket("Bob", "Alice", log_config=log_config)
3
4# Create an EPR socket for entanglement generation
5epr_socket = EPRSocket("Bob")
6
7with NetQASMConnection(epr_sockets=[epr_socket]) as conn:
8 # Create a qubit to teleport
9 q = Qubit(conn)
10
11 # Create EPR pairs
12 epr = epr_socket.create_keep()[0]
13
14 # Teleport
15 q.cnot(epr)
16 q.H()
17 m1 = q.measure()
18 m2 = epr.measure()
19
20# Send the correction information
21m1, m2 = int(m1), int(m2)
22
23socket.send_structured(StructuredMessage("Corrections", (m1, m2)))
Bob (Target Node)
# Create a classical socket
socket = Socket("Alice", "Bob")
# Create an EPR socket
epr_socket = EPRSocket("Alice")
with NetQASMConnection(epr_sockets=[epr_socket]) as conn:
# 1. Receive EPR qubit
epr = epr_socket.recv_keep()[0]
# 2. Flush to sync
conn.flush()
# 3. Get and apply the corrections
m1, m2 = socket.recv_structured().payload
if m2 == 1:
epr.X()
if m1 == 1:
epr.Z()
# 4. Final flush to sync
conn.flush()
Figure 2: Implementation of a teleport protocol using NetQASM SDK. The developer is responsible for manually orchestrating the EPR socket creation, classical message passing, and synchronization (flush).
III-B NetQASM SDK
To facilitate programming without writing raw assembly, the NetQASM SDK provides a Python interface for writing applications for specific network roles (e.g., “Alice” and “Bob”). A central abstraction in the SDK is the EPR Socket. By opening an EPR socket between two nodes, the programmer can request the generation of entanglement (e.g., creating Bell pairs) utilizing the underlying network link layer.
Figure 2 demonstrates the implementation of a teleport protocol using the NetQASM SDK, highlighting the steps required to transfer a qubit state from Alice to Bob.
III-C Programming Complexity
While the SDK abstracts the physical generation of entanglement, it does not provide high-level primitives for qubit transmission. Crucially, to transfer a quantum state from one node to another, the generation of an EPR pair is merely the first step. The programmer must manually implement the full quantum protocol communication ad hoc.
For example, to send a single qubit via teleportation [Bennett1993TeleportingChannels] using the SDK, the developer must explicitly:
@@ III-C Programming Complexity
    Create or request entanglement via the EPR Socket and wait for completion.
    • Alice: line 12.
    • Bob: line 9.
@@
IV NetQMPI
This section introduces NetQMPI, a high-level Python library designed to simplify the development of distributed quantum applications. Inspired by the standardized Message Passing Interface (MPI) used in classical HPC, NetQMPI adopts the Single Program Multiple Data (SPMD) paradigm, which enables developers to write a single, platform-independent source code that orchestrates quantum operations across multiple network nodes without manually managing the underlying physical entanglement resources.
IV-A Software stack
Figure 3 illustrates the architectural position of NetQMPI within the quantum network software ecosystem. NetQMPI acts as a middleware layer that sits directly on top of the NetQASM SDK.
Figure 3: The NetQMPI software stack. The library wraps the verbose NetQASM SDK primitives, providing a unified API for the application layer while relying on the standard NetQASM ISA to interface with backend simulators, such as SquidASM or SimulaQron, or real DQC clusters supporting NetQASM.
@@ IV-A Software stack
Because these primitives compile down to the standardized NetQASM ISA, NetQMPI is inherently backend-agnostic. This abstraction allows developers to write code that is completely decoupled from the underlying infrastructure, enabling the execution of the same source code on both simulated networks and physical quantum networks without modification.
@@
IV-B Construction of NetQMPI from NetQASM SDK
@@
The core orchestration is handled by the external.py module, which manages the multiprocess environment mapping. The network state and automated entanglement generation are encapsulated within the communicator package. Finally, the specific communication logic is organized under primitives, divided into p2p and collective modules, whose specific implementations will be detailed in Subsections IV-C and IV-D, respectively.
@@
IV-B1 Single Program Architecture
In the standard NetQASM SDK, developers must write $N$ separate source files (e.g., app_alice.py and app_bob.py), one for each node in the network. NetQMPI unifies this into a single file, following the SPMD paradigm.
To enable this, NetQMPI utilizes the abstractions of rank and size. For readers unfamiliar with HPC terminology, the rank is a unique integer identifier assigned to each node in the distributed network, ranging from 0 to size1size-1, while the size represents the total number of participating nodes in the communicator. These values are injected automatically by the library into the user’s code at runtime, allowing the developer to define the logic for all nodes in a single function signature.
@@ 6 COMM_WORLD = QMPICommunicator(rank, size, app_config)
Figure 5: An example of the rank and size variable injection with the global communicator creation.
When the user executes this script using the Command Line Interface (CLI), the cli.py module parses the -n argument (number of processes) and triggers the external.py dispatcher. This dispatcher spawns the simulation processes using the NetQASM SDK, passing the specific rank to each instance. For example, executing the code with three nodes results in:
@@
As illustrated in Figure 6, the execution flow is orchestrated by external.py. It parses the unified script, defines the network roles based on the requested size, and injects the corresponding rank information into the application’s entry point. Finally, it hands over the configuration to the underlying NetQASM SDK, which manages the physical spawning of the runtime processes (e.g., Alice, Bob, Charlie).
[…]
@@
IV-B2 Automated Network Initialization
The core abstraction that manages the logical connectivity between the distributed nodes is the QMPICommunicator class, defined in the communicator package.
Conceptually, a communicator acts as a virtualization layer that groups a set of logical processes connected through a logical network. This construct abstracts the underlying complexity of the infrastructure, where logical processes (identified by ranks) are mapped to physical resources (Quantum Processing Units) via a physical interconnection network. This relationship is visually described in Figure 7. Figure 7(a) depicts the physical reality where QPUs are connected via a Quantum Router, while Figure 7(b) illustrates how the software stack (OS/Middleware) maps the high-level Logical Ranks ($1\dots n$) to these physical resources transparently.
Figure 7: Logical abstraction for scheduling physical resources. (a) An example of physical resources interconnected in a physical quantum network. (b) The Communicator provides a logical layer where abstract Ranks are assigned to processes, and the OS/Middleware automatically handles the mapping to the available physical QPUs.
This approach mirrors the architecture of classical distributed computing, allowing developers to focus on the algorithmic logic rather than the physical topology. For a comprehensive definition of this abstraction model in distributed quantum computing, we refer the reader to the architectural framework presented in [Cardama2026NetQIR:Computing], specifically Figure 6.
From an implementation perspective, the communicator automates resource management that is currently handled manually in the NetQASM SDK. In NetQASM, establishing a network requires the user to instantiate a NetQASMConnection and define specific EPRSocket objects for every pair of communicating nodes, leading to an $O(N^{2})$ initialization complexity in fully connected topologies. NetQMPI resolves this by performing the following automated steps during initialization: